# CH2I2_UV_IR_CEI_SM

### v1

@@ Abstract
Femtosecond structural dynamics of diiodomethane ($\mathrm{CH_{2}I_{2}}$) triggered by ultraviolet (UV) photoabsorption at 290 nm and 330 nm are studied using time-resolved coincident Coulomb explosion imaging driven by a near-infrared probe pulse. We map the dominant single-photon process, the cleavage of the carbon-iodine bond producing rotationally excited $\mathrm{CH_{2}I}$ radical, identify the contributions of the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and molecular iodine formation channels, which are primarily driven by the absorption of more than one UV photon, and demonstrate the existence of a weak reaction pathway involving the formation of short-lived transient species resembling iso-$\mathrm{CH_{2}I-I}$ geometries with the slightly shorter I-I separation compared to the ground-state $\mathrm{CH_{2}I_{2}}$. These transient molecular configurations, which can be separated from the other channels by applying a set of conditions on the correlated momenta of three ionic fragments, are formed within approximately 100 fs after the initial photoexcitation and decay within the next 100 fs.
@@ I Introduction
Numerous studies have been published on the UV-induced photochemistry of $\mathrm{CH_{2}I_{2}}$, which have reported the primary cleavage of one of the C-I bonds and the formation of $\mathrm{CH_{2}I}$ and $\mathrm{I(^{2}P_{3/2})}$ or $\mathrm{I^{*}(^{2}P_{1/2})}$ photoproducts [20, 21, 22]. It has been concluded that $\mathrm{I(^{2}P_{3/2})}$ and $\mathrm{I^{*}(^{2}P_{1/2})}$ are obtained via photodissociation from the $\mathrm{1B_{1}}$ and $\mathrm{2B_{1}}$ states, respectively. Due to possible couplings between $\mathrm{2B_{1}}$ and higher-lying electronic states, $\mathrm{I(^{2}P_{3/2})}$ could also be produced via non-adiabatic dissociation from the $\mathrm{2B_{1}}$ state [22, 23]. Baughcum et al. [22] reported that the ratio of the yield of $\mathrm{I^{*}(^{2}P_{1/2})}$ and $\mathrm{I(^{2}P_{3/2})}$ decreases as the excitation wavelength increases, with the quantum yield for $\mathrm{I^{*}(^{2}P_{1/2})}$ production dropping from 0.46 at 248 nm to 0.25 at 308 nm. Using ions generated by a (2+1) resonance enhanced multiphoton ionization process, Xu et al.[23] measured the translational kinetic energy distributions of both $\mathrm{I(^{2}P_{3/2})}$ and $\mathrm{I^{*}(^{2}P_{1/2})}$ fragments in the wavelength range of 277-305 nm and concluded that the $\mathrm{CH_{2}I}$ co-fragment is produced with significant internal excitation, with approximately 80% of the total available energy being partitioned into the internal energy of the $\mathrm{CH_{2}I}$ fragment.
@@
Reid et al.[24, 25] later drew attention to the importance of isomerization of halogenated alkanes into iso-haloalkanes upon UV absorption as a major pathway leading to the production of molecular halogens. The photoisomerization of $\mathrm{CH_{2}I_{2}}$ to iso-$\mathrm{CH_{2}I_{2}}$ has been reported to be an extremely efficient process upon UV absorption in the solution phase and in cages[26, 27, 28]. Since it has been suggested that the dominant isomerization mechanisms in these studies are driven by the interaction with solvent molecules or cage, it is helpful to examine the process in isolated gas-phase molecules, free from such interactions. Borin et al.[29] reported observing photoisomerization of $\mathrm{CH_{2}I_{2}}$ in the gas phase upon 330 nm UV excitation by probing it with femtosecond transient absorption. The delayed rise in the transient absorption of $\mathrm{CH_{2}I_{2}}$ at 380 and 612 nm, approximately 35-90 fs after UV excitation, was attributed to the creation of a short-lived iso-$\mathrm{CH_{2}I_{2}}$. However, a recent ultrafast electron diffraction (UED) study, where Liu et al.[30] directly mapped the C-I cleavage upon 266 nm excitation, quantified the resulting increase of the C-I and I-I separations, and observed signatures of the rotation of the $\mathrm{CH_{2}I}$ radical after the C-I bond cleavage, did not mention any observation of iso-$\mathrm{CH_{2}I_{2}}$ formation.
@@
Here, we employ Coulomb Explosion Imaging (CEI), a powerful method for determining the geometric structure of gas-phase molecules [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], to investigate the structural dynamics of $\mathrm{CH_{2}I_{2}}$ upon UV photoabsorption. CEI has been shown to be a useful tool in identifying molecular isomers [35] and conformers [36] of polyhalogenated alkanes and other organic molecules [42], and even imaging the complete structure of halogenated alkanes [38, 39] and ring molecules [37, 40, 42]. When used as a time-resolved method in a pump-probe scheme, laser-induced CEI can measure wave-packets dynamics[44, 45, 46], visualize vibrational motions[47, 48, 46], and image photo-dissociation[49, 50, 51, 52] with high temporal resolution[53]. In particular, CEI provides multidimensional data crucial for identifying any transient structural changes in the molecule upon UV excitation, often with the help of simple classical modeling and computations.
@@
In this work, we investigate the dynamics of diiodomethane upon UV excitation using ionization and Coulomb explosion by an intense strong-field near-infrared (NIR) probe pulse. The primary objective of this time-resolved study is to explore possible signatures of the intramolecular photoisomerization process in $\mathrm{CH_{2}I_{2}}$, guided by the findings reported by Borin et al. [29]. We analyze the delay-dependent angular correlations between the momenta of the ions, the total kinetic energy release (KER), and the kinetic energies (KE) of individual ionic fragments to differentiate the possible photoisomerization channel from the direct two-body break-up channel that leads to $\mathrm{CH_{2}I+I}$ (or $\mathrm{I^{*}}$) products. We also identify the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and the I2 formation channels, which primarily occur after the absorption of multiple UV photons. A schematic of the pump-probe experiment and the different reaction pathways after UV photoexcitation is shown in Fig. 1.
@@ II.1 Experimental
Figure 2: Schematic of the pump-probe setup and the coincident ion momentum imaging system. The diagram depicts the propagation of the ultraviolet (UV) pump and near-infrared (NIR) probe pulses into the experimental chamber. Both pulses are colinearly directed and focused onto a cold supersonic molecular jet containing diiodomethane. The setup is integrated with a double-sided velocity map imaging (VMI) spectrometer, operated in multi-ion coincidence mode. Ions are detected in coincidence using a time and position-sensitive detector comprising a microchannel plate (MCP) and delay line anode assembly. It should be noted that in this particular experiment, the electron detector is not utilized.
Figure 2 illustrates the experimental setup for the UV-pump and NIR-probe CEI experiment. The femtosecond laser system comprises of a 5-mJ, 3-kHz Coherent Legend Elite DUO Ti:Sapphire laser coupled to a Light Conversion TOPAS Prime optical parametric amplifier (OPA). The output of the Ti:Sapphire laser is split evenly into two independently compressed beam paths: one driving the OPA, and the other providing near-Fourier-transform-limited NIR pulses with a full width at half maximum intensity (FWHM) duration of 26 fs and a bandwidth of 60 nm at a central wavelength of 810 nm. In this experiment, the OPA was set to produce pulses at a central wavelength of 290 nm and 330 nm, with a 5-nm bandwidth, pulse duration of 70 fs (FWHM), and pulse energy up to 15 ${\mu}$J. The UV and NIR pulses were recombined using a 45-degree recombination mirror that transmitted the NIR pulses and reflected the UV pulses, and the combined beam was directed collinearly into the vacuum chamber through a 1 mm calcium fluoride $\mathrm{(CaF_{2})}$ window. The delay between NIR and UV pulses was adjusted via a computer-controlled optical delay stage in the NIR arm. Inside the vacuum chamber, the laser pulses are focused using a normal-incidence spherical mirror with 75-mm focal length and UV enhanced aluminium coating.
@@ II.2 Coulomb explosion simulation
To guide the search for a possible isomer and identify the different reaction channels, we performed classical Coulomb explosion simulations for the molecular geometries corresponding to the $\mathrm{CH_{2}I_{2}}$ equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$, and the C-I bond cleavage pathway including rotation of the $\mathrm{CH_{2}I}$ intermediate. The simulations are based on solving Newton’s equations of motion for three-point charges located at the positions of the carbon and iodine atoms. To match the experimental observables, the $\mathrm{CH_{2}}$ fragment was kept intact. It is treated as a point particle in the simulations. Furthermore, the ionization and fragmentation process induced by the probe pulse was assumed to be instantaneous (i.e., instantaneously breaking both C-I bonds and resulting in a charge q=1 on the carbon atom and q=2 on each of the two iodine atoms), and the repulsion between the three fragments was treated as purely Coulombic. From prior experience, the calculations based on these assumptions typically overestimate the resulting KER but reproduce the experimentally observed momentum correlations [57, 58, 59, 60, 35, 36, 38, 42].
The optimized structures of the molecule in the equilibrium geometry of the neutral electronic ground state and the isomer geometry were taken from Borin et al. [29]. For each case, an ensemble of 5000 geometries was generated by randomly varying the initial atomic positions within a radius of 0.1 Å around the equilibrium geometries and the initial velocities by up to $\mathrm{2x10^{-4}}$ a.u., respectively. These parameters were chosen empirically in order to approximately match the width of the fragment kinetic energy and momentum distributions observed in the experiment. In the simulation for the dissociation channel, a rotational period of 300 fs (obtained experimentally) was used for the $\mathrm{CH_{2}I}$ intermediate, and the translational KEs used for the ion pairs were obtained from the experimental asymptotic KER for the respective channels.
@@ III Results and discussions
While the main objective of this work is to investigate the transient photoisomerization of diiodomethane ($\mathrm{CH_{2}I}$-$\mathrm{I}$) after UV photoabsorption, the majority of the photoexcited molecules undergo direct dissociation following C-I bond cleavage, and the first task is therefore to identify the signatures of these competing reaction pathways. The intense NIR probe pulse ionizes both, the unpumped and the photoexcited molecules, resulting in multiple channels with different final charge states and different ionic fragments (see Figs. S1 - S4 of the Supplementary Material (SM) for ion time-of-flight (ToF) mass spectra and multi-ion coincidence spectra). The following discussion focuses on the analysis of the $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ coincidence channel because, as discussed in detail in[61], this breakup channel contains minimal contribution from sequential fragmentation, which is prominent for channels with lower total charge states (see Fig. S5 and Fig. S6 of the SM), where it complicates the analysis. In Fig. 3a, the ion yield of this coincident channel is plotted as a function of the KER and the pump-probe delay.
@@
Figure 3: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of pump-probe delay and KER. Positive delays correspond to the NIR probe pulse arriving after the UV pump pulse. A step size of 15 fs and 100 fs was used for delays up to and beyond 800 fs, respectively. (b) KER spectra of the same coincidence channel recorded at a fixed delay of 13.5 ps. The data shown here were recorded at a pump wavelength of 290 nm. The equivalent plots for this and all the following figures with the data recorded at 330 nm are shown in Section II of the SM.
@@
Figure 4: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of KER and angle between the momentum vectors of the two iodine dications for a fixed pump-probe delay of 13.5 ps. Four distinct regions are marked by colored ovals corresponding to different contributions to the coincidence ion yield: Black: Coulomb explosion of bound molecules in or near the equilibrium geometry. Red: C-I cleavage and dissociation to $\mathrm{CH_{2}I}$ + $\mathrm{I}$. Green: Molecular iodine ($\mathrm{I_{2}}$) formation after UV absorption. Blue: UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. (b) Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the same $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ channel observed in the experiment, for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [29], and for the three dissociation processes that are identified and marked in (a).
@@
Three distinct features are visible in this plot: (i) a horizontal band in the KER range between 30 and 40 eV, which corresponds to the Coulomb explosion of bound (i.e. non-dissociating) molecules that may or may not have absorbed a UV photon and that were five-fold ionized by the probe pulse; (ii) a curved feature whose KER decreases with increasing pump-probe delay and which emerges from the horizontal band near time-zero, reaching a KER of approximately 15-18 eV at a delay of 2 ps and 9-14 eV at a delay of 13.5 ps (see Fig. 3b); and (iii) another curved feature whose KER decreases even faster than that of feature (ii), reaching a KER of approximately 8-10 eV at a delay of 2 ps and 1-5 eV at a delay of 13.5 ps, at which point the KER no longer changes with delay. Features (ii) and (iii) originate from the molecules that dissociated upon UV photoabsorption, with feature (ii) corresponding to two-body dissociations into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ or $\mathrm{CH_{2}}$ and $\mathrm{I_{2}}$, while feature (iii) corresponds to direct three-body dissociation by the UV pulse into $\mathrm{CH_{2}}$, I and I. While these assignments can be made with good confidence based on the asymptotic KERs of both channels, further insights and confirmation are obtained when inspecting the angular correlation between the momentum vectors of the ionic fragments[61].
@@
Figure 4a shows the coincidence ion yield at the asymptotic delay of 13.5 ps as a function of the KER and the angle between the momentum vectors of two iodine dications. The most intense feature at a KER of 30 - 40 eV (with a smaller side-peak between 20 - 30 eV) and relatively well defined ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle centered around 150 - marked by the black ellipse in Fig. 4a - stems from the Coulomb explosion of bound molecules in or near their equilibrium geometry, as confirmed by the data taken without the UV pulse present (see Fig. S6 in the SM). At a KER of approximately 15 eV (red ellipse), an angularly broad feature spanning 60-180 degrees in the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle can be attributed to C-I cleavage and dissociation into $\mathrm{CH_{2}I}$ + $\mathrm{I}$ induced by a single-UV-photon absorption. The broad spread in angle results from high rotational excitation of the $\mathrm{CH_{2}I}$ fragment due to the torque imparted from the C-I bond cleavage, as also observed in other dihalomethanes [62, 63, 64, 65, 66, 67].
@@
Two additional features, each of them with approximately half the number of events as in the red ellipse discussed above, are prominent in Fig. 4. First, a rather localized spot peaked at a KER slightly below 20 eV and an ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle close to 180° (green ellipse) can be uniquely attributed to $\mathrm{I_{2}}$ formation after UV excitation (i.e., dissociation into $\mathrm{CH_{2}}$ + $\mathrm{I_{2}}$) since both the KER and the back-to-back emission of the two iodine ions is consistent with a Coulomb explosion of $\mathrm{I_{2}}$. Second, a clearly separated contribution at low KER (blue ellipse) is attributed to the UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. Both of these features result mainly from multi-photon excitation by the UV pulse. This assignment is based on (i) the energy required to trigger three-body dissociation (4.8 - 5 eV [68, 69]), (ii) earlier experimental work that reported rather low quantum yield of $\mathrm{I_{2}}$ elimination upon single-photon excitation [68], and (iii) on the observed dependence on the UV power. A detailed analysis of the UV power dependence of the different channels for pump wavelengths of 290 nm and 330 nm is presented in Fig. S12 of the SM.
@@
The results of the classical Coulomb explosion simulations for the neutral ground-state geometry and the iso-$\mathrm{CH_{2}I_{2}}$ equilibrium geometry are shown in Fig. 4b, along with the simulation results for the $\mathrm{CH_{2}I-I}$, $\mathrm{CH_{2}-I_{2}}$ and $\mathrm{CH_{2}-I-I}$ dissociation channels at 13.5 ps pump-probe delay. Apart from overestimating the measured KER, most likely due to the assumption of a purely Coulombic potential, the simulated results qualitatively match the experimental observations for all channels, validating the use of these simulations in identifying the observables corresponding to the possible formation of the isomer.
@@
Although the simulations show a subtle difference in the KER of the equilibrium geometry and the isomer, this distinction can be expected to be much less pronounced in the experiment, where a significant spread in the KER is observed even when no UV pulse is present. However, the simulations also predict that the Coulomb explosion of the isomer will lead to some ion yield at larger angles between the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) momenta than those realized by the Coulomb explosion of the equilibrium geometry. In the experiment, essentially no events are observed with ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angles larger than 160° when no UV pulse is present (see Fig. S6 in the SM), but a transient signal with high KER is observed in this region for pump probe delays between 0 and 250 fs (see Supplementary Movie 1). By selecting only those events in the experimental data where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160°, we can thus almost completely suppress the contribution from molecules in the equilibrium geometry.

### v2

# nocite]CH2I2_UV_IR_CEI_SM

@@ Abstract
Femtosecond structural dynamics of diiodomethane ($\mathrm{CH_{2}I_{2}}$) triggered by ultraviolet (UV) photoabsorption at 290 nm and 330 nm are studied using time-resolved coincident Coulomb explosion imaging driven by a near-infrared probe pulse. We map the dominant single-photon process, the cleavage of the carbon-iodine bond producing rotationally excited $\mathrm{CH_{2}I}$ radical, identify the contributions of the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and molecular iodine formation channels, which are primarily driven by the absorption of more than one UV photon, and demonstrate the existence of a weak reaction pathway involving the formation of short-lived transient species resembling iso-CH2I2-like geometries with a slightly shorter I–I separation compared to the ground-state $\mathrm{CH_{2}I_{2}}$. These transient molecular configurations, which can be separated from the other channels by applying a set of conditions on the correlated momenta of three ionic fragments, are formed within approximately 100 fs after the initial photoexcitation and decay within the next 100 fs.
@@ I Introduction
Numerous studies have been published on the UV-induced photochemistry of $\mathrm{CH_{2}I_{2}}$, which have reported the primary cleavage of one of the C–I bonds and the formation of $\mathrm{CH_{2}I}$ and $\mathrm{I(^{2}P_{3/2})}$ or $\mathrm{I^{*}(^{2}P_{1/2})}$ photoproducts [28, 34, 5]. Using ions generated by a (2+1) resonance enhanced multiphoton ionization process, Xu et al.[70] measured the translational kinetic energy distributions of both $\mathrm{I(^{2}P_{3/2})}$ and $\mathrm{I^{*}(^{2}P_{1/2})}$ fragments in the wavelength range of 277–305 nm and concluded that the $\mathrm{CH_{2}I}$ co-fragment is produced with significant internal excitation, with approximately 80% of the total available energy being partitioned into the internal energy of the $\mathrm{CH_{2}I}$ fragment.
@@
Reid et al.[26, 54] later drew attention to the importance of isomerization of halogenated alkanes into iso-haloalkanes upon UV absorption as a major pathway leading to the production of molecular halogens. The photoisomerization of $\mathrm{CH_{2}I_{2}}$ to iso-$\mathrm{CH_{2}I_{2}}$ has been reported to be an extremely efficient process upon UV absorption in the solution phase and in cages[63, 49, 31]. Since it has been suggested that the dominant isomerization mechanisms in these studies are driven by the interaction with solvent molecules or cage, it is helpful to examine the process in isolated gas-phase molecules, free from such interactions. Borin et al.[10] reported observing photoisomerization of $\mathrm{CH_{2}I_{2}}$ in the gas phase upon 330 nm UV excitation by probing it with femtosecond transient absorption. The delayed rise in the transient absorption of $\mathrm{CH_{2}I_{2}}$ at 380 and 612 nm, approximately 35–90 fs after UV excitation, was attributed to the creation of a short-lived iso-$\mathrm{CH_{2}I_{2}}$. A recent ultrafast electron diffraction (UED) study, where Liu et al.[45] directly mapped the C–I cleavage upon 266 nm excitation, quantified the resulting increase of the C–I and I–I separations, and observed signatures of the rotation of the $\mathrm{CH_{2}I}$ radical after the C–I bond cleavage, but did not mention any observation of iso-$\mathrm{CH_{2}I_{2}}$ formation.
@@
Here, we employ Coulomb Explosion Imaging (CEI), a powerful method for determining the geometric structure of gas-phase molecules [65, 29, 13, 14, 2, 50, 9, 6, 42, 37, 68, 36, 24, 66, 43], to investigate the structural dynamics of $\mathrm{CH_{2}I_{2}}$ upon UV photoabsorption, specifically at 290 nm and 330 nm. Excitation at both 290 nm (4.28 eV) and 330 nm (3.76 eV) can access several excited states that are adiabatically leading to CH2I+I(2P3/2) dissociation. Excitation at 290 nm can also access states leading to CH2I+I(2P1/2) dissociation[64]. CEI has been shown to be a useful tool in identifying molecular isomers [2] and conformers [50] of polyhalogenated alkanes and other organic molecules [36], and even imaging the complete structure of halogenated alkanes [6, 42] and ring molecules [9, 37, 36]. When used as a time-resolved method in a pump-probe scheme, laser-induced CEI can measure wave-packets dynamics[61, 38, 35], visualize vibrational motions[62, 16, 35], and image photo-dissociation[8, 4, 71, 7] with high temporal resolution[39]. In particular, CEI provides multidimensional data crucial for identifying any transient structural changes in the molecule upon UV excitation, often with the help of simple classical modeling and computations.
@@
In this work, we investigate the dynamics of diiodomethane upon UV excitation using ionization and Coulomb explosion by an intense strong-field near-infrared (NIR) probe pulse. The primary objective of this time-resolved study is to explore possible signatures of the intramolecular photoisomerization process in $\mathrm{CH_{2}I_{2}}$, guided by the findings reported by Borin et al. [10]. We analyze the delay-dependent angular correlations between the momenta of the ions, the total kinetic energy release (KER), and the kinetic energies (KE) of individual ionic fragments to differentiate the possible photoisomerization channel from the direct two-body breakup channel that leads to $\mathrm{CH_{2}I+I}$ (or $\mathrm{I^{*}}$) products. We also identify the three-body ($\mathrm{CH_{2}+I+I}$) dissociation and the I2 formation channels, which primarily occur after the absorption of multiple UV photons. A schematic of the pump-probe experiment and the different reaction pathways after UV photoexcitation is shown in Fig. 1.
@@ II.1 Experimental
Figure 2: Schematic of the pump-probe setup and the coincident ion momentum imaging system. The diagram depicts the propagation of the ultraviolet (UV) pump and near-infrared (NIR) probe pulses into the experimental chamber. Both pulses are collinearly directed and focused onto a cold supersonic molecular jet containing diiodomethane. The setup is integrated with a double-sided velocity map imaging (VMI) spectrometer, operated in multi-ion coincidence mode. Ions are detected in coincidence using a time- and position-sensitive detector comprising a microchannel plate (MCP) and delay line anode assembly. It should be noted that in this particular experiment, the electron detector is not utilized.
Figure 2 illustrates the experimental setup for the UV-pump and NIR-probe CEI experiment. The femtosecond laser system comprises a 5-mJ, 3-kHz Coherent Legend Elite DUO Ti:Sapphire laser coupled to a Light Conversion TOPAS Prime optical parametric amplifier (OPA). The output of the Ti:Sapphire laser is split evenly into two independently compressed beam paths: one driving the OPA, and the other providing near-Fourier-transform-limited NIR pulses with a full width at half maximum intensity (FWHM) duration of 26 fs and a bandwidth of 60 nm at a central wavelength of 810 nm. In this experiment, the OPA was set to produce pulses at a central wavelength of 290 nm and 330 nm, with a 5-nm bandwidth, pulse duration of 70 fs (FWHM), and pulse energy up to 15 $\mu$}J. The UV and NIR pulses were recombined using a 45-degree recombination mirror that transmitted the NIR pulses and reflected the UV pulses, and the combined beam was directed collinearly into the vacuum chamber through a 1-mm calcium fluoride $\mathrm{(CaF_{2})}$ window. The delay between NIR and UV pulses was adjusted via a computer-controlled optical delay stage in the NIR arm. Inside the vacuum chamber, the laser pulses are focused using a normal-incidence spherical mirror with 75-mm focal length and UV-enhanced aluminum coating.
@@ II.2 Coulomb explosion simulation
To guide the search for a possible isomer and identify the different reaction channels, we performed classical Coulomb explosion simulations for the molecular geometries corresponding to the $\mathrm{CH_{2}I_{2}}$ equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$, and the C–I bond cleavage pathway including rotation of the $\mathrm{CH_{2}I}$ intermediate. The simulations are based on solving Newton’s equations of motion for three point charges located at the positions of the carbon and iodine atoms. To match the experimental observables, the $\mathrm{CH_{2}}$ fragment was kept intact. It is treated as a point particle in the simulations. Furthermore, the ionization and fragmentation process induced by the probe pulse was assumed to be instantaneous (i.e., instantaneously breaking both C–I bonds and resulting in a charge q=1 on the carbon atom and q=2 on each of the two iodine atoms), and the repulsion between the three fragments was treated as purely Coulombic. From prior experience, the calculations based on these assumptions typically overestimate the resulting KER but reproduce the experimentally observed momentum correlations [41, 40, 44, 15, 2, 50, 6, 36].
The optimized structures of the molecule in the equilibrium geometry of the neutral electronic ground state and the isomer geometry were taken from Borin et al. [10]. For each case, an ensemble of 5000 geometries was generated by randomly varying the initial atomic positions within a radius of 0.1 Å around the equilibrium geometries and the initial velocities by up to $\mathrm{2x10^{-4}}$ a.u., respectively. These parameters were chosen empirically in order to approximately match the width of the fragment kinetic energy and momentum distributions observed in the experiment. In the simulation for the dissociation channel, a rotational period of $\sim$300 fs (obtained experimentally for the two-body dissociation into $\mathrm{CH_{2}I}$ + I following UV absorption and C–I cleavage, see Fig. S11) was used for the $\mathrm{CH_{2}I}$ intermediate, and the translational KEs used for the ion pairs were obtained from the experimental asymptotic KER for the respective channels.
@@ III Results and discussions
Figure 3: Newton plots of the (a) $\mathrm{CH_{2}^{+}+I^{+}+I^{+}}$, (b) $\mathrm{CH_{2}^{+}+I^{2+}+I^{+}}$, and (c) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ coincidence channels for ionization by the NIR pulse alone. The momenta are normalized to the magnitude of the $\mathrm{CH_{2}^{+}}$ fragment, and the reference frame is chosen such that the $\mathrm{CH_{2}^{+}}$ momentum is aligned along the $x$-axis, with the xyxy plane defined by the first detected iodine ion in each channel. The circular features in (a) and (b) are clear indications of sequential breakup.
[…]
@@
Figure 4: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of pump-probe delay and KER. Positive delays correspond to the NIR probe pulse arriving after the UV pump pulse. A step size of 15 fs and 100 fs was used for delays up to and beyond 800 fs, respectively. (b) KER spectra of the same coincidence channel recorded at a fixed delay of 13.5 ps. The data shown here were recorded at a pump wavelength of 290 nm. The equivalent plots for this and all the following figures with the data recorded at 330 nm are shown in Section II of the SM. The vertical dashed line at 300 fs in panel (a) indicates the experimentally obtained rotational period of the $\mathrm{CH_{2}I}$ fragment.
@@
Figure 5: (a) $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ ion coincidence yield as a function of KER and angle between the momentum vectors of the two iodine dications for a fixed pump-probe delay of 13.5 ps. Four distinct regions are marked by colored ovals corresponding to different contributions to the coincidence ion yield: Black: Coulomb explosion of bound molecules in or near the equilibrium geometry. Red: C–I cleavage and dissociation to $\mathrm{CH_{2}I}$ + $\mathrm{I}$. Green: Molecular iodine ($\mathrm{I_{2}}$) formation after UV absorption. Blue: UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. (b) Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the same $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ channel observed in the experiment, for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [10], and for the three dissociation processes that are identified and marked in (a). The white contour line in (b) marks the region containing 90% of the simulated iso-$\mathrm{CH_{2}I_{2}}$ events, showing its partial overlap with the simulated events from the ground-state equilibrium geometries.
@@
Three distinct features are visible in this plot: (i) a horizontal band in the KER range between 30 and 40 eV, which corresponds to the Coulomb explosion of bound (i.e. non-dissociating) molecules that may or may not have absorbed a UV photon and that were five-fold ionized by the probe pulse; (ii) a curved feature whose KER decreases with increasing pump-probe delay and which emerges from the horizontal band near time-zero, reaching a KER of approximately 15–18 eV at a delay of 2 ps and 9–14 eV at a delay of 13.5 ps (see Fig. 4(b)); and (iii) another curved feature whose KER decreases even faster than that of feature (ii), reaching a KER of approximately 8–10 eV at a delay of 2 ps and 1–5 eV at a delay of 13.5 ps, at which point the KER no longer changes with delay. Features (ii) and (iii) originate from the molecules that dissociated upon UV photoabsorption, with feature (ii) corresponding to two-body dissociations into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ or $\mathrm{CH_{2}}$ and $\mathrm{I_{2}}$, while feature (iii) corresponds to direct three-body dissociation by the UV pulse into $\mathrm{CH_{2}}$, I and I. While these assignments can be made with good confidence based on the asymptotic KERs of both channels, further insights and confirmation are obtained when inspecting the angular correlation between the momentum vectors of the ionic fragments[25].
@@
Figure 5(a) shows the coincidence ion yield at the asymptotic delay of 13.5 ps as a function of the KER and the angle between the momentum vectors of two iodine dications. The most intense feature at a KER of 30–40 eV (with a smaller side-peak between 20–30 eV) and relatively well defined ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle centered around 150 (marked by the black ellipse in Fig. 5(a)) stems from the Coulomb explosion of bound molecules in or near their equilibrium geometry, as confirmed by the data taken without the UV pulse present (see Fig. S4 in the SM). At a KER of approximately 15 eV (red ellipse), an angularly broad feature spanning 60–180 degrees in the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle can be attributed to C–I cleavage and dissociation into $\mathrm{CH_{2}I}$ + $\mathrm{I}$ induced by a single-UV-photon absorption. The broad spread in angle results from high rotational excitation of the $\mathrm{CH_{2}I}$ fragment due to the torque imparted from the C–I bond cleavage, as also observed in other dihalomethanes [48, 3, 47, 32, 53, 67].
@@
Two additional features, each of them with approximately half the number of events as in the red ellipse discussed above, are prominent in Fig. 5. First, a rather localized spot peaked at a KER slightly below 20 eV and an ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle close to 180° (green ellipse) can be uniquely attributed to $\mathrm{I_{2}}$ formation after UV excitation (i.e., dissociation into $\mathrm{CH_{2}}$ + $\mathrm{I_{2}}$) since both the KER and the back-to-back emission of the two iodine ions are consistent with a Coulomb explosion of $\mathrm{I_{2}}$. Second, a clearly separated contribution at low KER (blue ellipse) is attributed to the UV-induced three-body dissociation into $\mathrm{CH_{2}}$ + I + I. Both of these features result mainly from multi-photon excitation by the UV pulse. This assignment is based on (i) the energy required to trigger three-body dissociation (4.8 - 5 eV [12, 64]), (ii) earlier experimental work that reported rather low quantum yield of $\mathrm{I_{2}}$ elimination upon single-photon excitation [12], and (iii) on the observed dependence on the UV power. The log–log analysis, as shown in Fig. S15, reveals that the C–I dissociation yield scales linearly with pump intensity (slope $\approx 1$), consistent with a one-photon absorption process. In contrast, the molecular I2 formation and the CH2 + I + I three-body fragmentation scale nonlinearly (slope $\approx 2$), confirming their two-photon character. A detailed analysis of the UV power dependence of the different channels for pump wavelengths of 290 nm and 330 nm is presented in Sec. IV of the SM.
@@
The results of the classical Coulomb explosion simulations for the neutral ground-state geometry and the iso-$\mathrm{CH_{2}I_{2}}$ equilibrium geometry are shown in Fig. 5(b), along with the simulation results for the $\mathrm{CH_{2}I}$ + I, $\mathrm{CH_{2}}$ + I2 and $\mathrm{CH_{2}}$ + I + I dissociation channels at 13.5 ps pump-probe delay. Apart from overestimating the measured KER, most likely due to the assumption of a purely Coulombic potential, the simulated results qualitatively match the experimental observations for all channels, validating the use of these simulations in identifying the observables corresponding to the possible formation of the isomer.
@@
Although the simulations show a subtle difference in the KER of the equilibrium geometry and the isomer, this distinction can be expected to be much less pronounced in the experiment, where a significant spread in the KER is observed even when no UV pulse is present. However, the simulations also predict that the Coulomb explosion of the isomer will lead to some ion yield at larger angles between the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) momenta than those realized by the Coulomb explosion of the equilibrium geometry. In the experiment, essentially no events are observed with ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angles larger than 160 when no UV pulse is present (see Fig. S4 in the SM), but a transient signal with high KER is observed in this region for pump probe delays between 0 and 250 fs, see movie in Fig. S16 (Multimedia available online). By selecting only those events in the experimental data where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160, we can thus almost completely suppress the contribution from molecules in the equilibrium geometry.